As sales manager, Joe Batista was given the following static budget report for selling expenses in the Clothing Department of Soria Company for the month of October.

SORIA COMPANY Clothing Department Budget Report For the Month Ended October 31, 2020 Difference Budget Actual Favorable Unfavorable Neither Favorable nor Unfavorable Sales in units 8,500 10,000 1,500 Favorable Variable expenses Sales commissions $2,380 $2,600 $220 Unfavorable Advertising expense 850 900 50 Unfavorable Travel expense 3,910 3,500 410 Favorable Free samples given out 1,615 1,100 515 Favorable Total variable 8,755 8,100 655 Favorable Fixed expenses Rent 1,700 1,700 –0– Neither Favorable nor Unfavorable Sales salaries 1,000 1,000 –0– Neither Favorable nor Unfavorable Office salaries 600 600 –0– Neither Favorable nor Unfavorable Depreciation—autos (sales staff) 500 500 –0– Neither Favorable nor Unfavorable Total fixed 3,800 3,800 –0– Neither Favorable nor Unfavorable Total expenses $12,555 $11,900 $655 Favorable As a result of this budget report, Joe was called into the president’s office and congratulated on his fine sales performance. He was reprimanded, however, for allowing his costs to get out of control. Joe knew something was wrong with the performance report that he had been given. However, he was not sure what to do, and comes to you for advice. Prepare a budget report based on flexible budget data to help Joe.

(List variable costs before fixed costs.) SORIA COMPANY Selling Expense Flexible Budget Report Clothing Department For the Month Ended October 31, 2020 Difference Budget Actual Favorable Unfavorable Neither Favorable nor Unfavorable $ $ $ $ $ $
Answer #1
SORIA COMPANY
Selling Expense Flexible Budget Report
For the Month Ended October 31, 2020
Budget Actual Difference
Sales in units 10000 10000
Variable costs:
Sales commissions 2800 2600 200 Favorable
Advertising expense 1000 900 100 Favorable
Travel expense 4600 3500 1100 Favorable
Free samples given 1900 1100 800 Favorable
Total variable 10300 8100 2200 Favorable
Fixed costs:
Rent 1700 1700 0 Neither Favorable nor Unfavorable
Sales salaries 1000 1000 0 Neither Favorable nor Unfavorable
Office salaries 600 600 0 Neither Favorable nor Unfavorable
Depreciation—autos 500 500 0 Neither Favorable nor Unfavorable
Total fixed 3800 3800 0 Neither Favorable nor Unfavorable
Total costs 14100 11900 2200 Favorable
Workings:
Budget
Sales commissions 2800 =2380/8500*10000
Advertising expense 1000 =850/8500*10000
Travel expense 4600 =3910/8500*10000
Free samples given 1900 =1615/8500*10000